OpenStudy (anonymous):

Circle A has an equation of (x + 8)2 + y2 = 16. Where is the center of the circle located?

OpenStudy (anonymous):

(-8,0)

OpenStudy (callisto):

Equation in the centre-radius form (x -h)^2 + (y-k)^2 = r^2 where (h,k) is the centre of the circle, r is the radius So, compare this with your question (x + 8)2 + y2 = 16 h = -8 and k=0 This would give you centre = (-8,0)

OpenStudy (lgbasallote):

the equation of a circle is (x-h)^2 + (y-k)^2 = r^2 where (h,k) is the center...rephrasing... (x- (-8)) + (y-0)^2 = 16 so h is -8 k is 0